OpenStudy (anonymous):

A solution of HCl is 0.300M determine what mass of acid has been dissolved in 150 mL of solution

OpenStudy (abmon98):

Number of Moles=Concentration(mol/dm^3)*Volume(dm^3) Change 150 ml unit into dm^3 by dividing by 1000 150/1000*0.300=0.045 moles Number of Moles=Mass(g)/Molar Mass(g/mol) Use your periodic table to find each element's atomic weight of HCL H:1 Cl:35.5 Molar Mass:(35.5*1)+(1*1)=36.5 grams Arrange Number of Moles=Mass(g)/Molar Mass(g/mol) to Number of Moles*Molar Mass(g/mol)=Mass(g) 36.5*0.045=1.6425 grams